### v3

@@ 3.2 MDP Formulation
We formulate stock and cryptocurrency trading tasks as Markov Decision Processes (MDPs) 10.
@@
  • Action $\mathbf{a_{t}}\in\mathbb{R}^{K}$ represents trading actions at time $t$, such that $\mathbf{h}_{t+1}=\max((\mathbf{h_{t}}+\mathbf{a_{t}}),0)$. An entry $a_{t}^{i}>0,i=1,\ldots,K$ indicates buying $a_{t}^{i}$ shares of assets $i$, while $a_{t}^{i}<0$ indicates selling and $a_{t}^{i}=0$ indicates holding the current position. Each entry of $\mathbf{h_{t+1}}$ is nonnegative, which means the agent cannot sell more shares than it currently holds.
@@
  • Reward $r_{t+1}=R(\mathbf{s_{t}},\mathbf{a_{t}},\mathbf{s_{t+1}})$ is an incentive signal to motivate the trading agent to perform action $\mathbf{a_{t}}$ at state $\mathbf{s_{t}}$. We use $r_{t+1}$ instead of $r_{t}$ to reflect the delayed reward in financial markets. It emphasizes that the reward $r_{t+1}$ is determined at the next state $\mathbf{s}_{t+1}$ after taking action $\mathbf{a}_{t}$. The reward can be defined as the change in the total asset value, i.e., $R(\mathbf{s_{t}},\mathbf{a_{t}},\mathbf{s_{t+1}})=v_{t+1}-v_{t}$, where $v_{t}=b_{t}+\mathbf{p_{t}}^{T}\mathbf{h_{t}}$ is the total asset value at time $t$. The reward can also be penalized for high-risk signals.
@@
In the state $\mathbf{s}_{t}$, the holding positions $\mathbf{h}_{t}$ are typically constrained to be nonnegative in long-only trading settings, where the agent is not allowed to short assets. The feature vector $\mathbf{f}_{t}$ can include market indicators (e.g., moving average convergence divergence), factors learned by neural networks, and LLM-generated signals from financial news or regulatory filings. The action $\mathbf{a}_{t}$ can be discrete or continuous, depending on the task. It may support long-only or long-short strategies. The reward function $R$ can use the Sharpe ratio in addition to the change in asset value. It can be further adjusted by risk or market sentiment to better reflect real-world trading scenarios.
For example, a trading task is to develop a trading agent for all 30 constituent stocks in the Dow Jones Index with an initial investment of $\mathdollar 1$ million. In this setting, the number of assets is $K=30$. The state includes the account balance with $b_{0}=100,000$, prices of the 30 stocks, holding positions of the 30 stocks, and $I=4$ technical indicators for each stock. The state space has a dimension $K(I+2)+1=181$. The action space has dimension $K=30$. An entry, for example, $a_{t}^{1}=5$ means buying $5$ shares of the first stock in the portfolio. The reward function is the change in the total asset value, with $v_{0}=100,000$. Taking the first stock as an example, assume at time $t$, the account balance is $b_{t}=100,000$, the price of the first stock is $p_{t}^{1}=200$, and the shares of the first stock are $h_{t}^{1}=10$. The agent buys 5 shares of the first stock ($a_{t}^{1}=5$) and the price increases to $p_{t+1}^{1}=201$ at time $t+1$. Assuming holding all other 29 stocks, the state is updated where $b_{t+1}=99,000$, $p_{t+1}^{1}=201$, and $h_{t+1}=15$. The return $r_{t}=v_{t+1}-v_{t}=(b_{t+1}+\mathbf{p}_{t+1}^{T}\mathbf{h}_{t+1})-(b_{t}+\mathbf{p}_{t}^{T}\mathbf{h}_{t})=b_{t+1}-b_{t}+p_{t+1}^{1}h_{t+1}^{1}-p_{t}^{1}h_{t}^{1}=-1000+201\cdot 15-200\cdot 10=15$. In this example, the agent only operates on the first stock.
@@ 3.3 Policy Gradient Estimate and Its Parallelism
In the FinRL tasks, the policy $\pi_{\theta}$ is typically a neural network, where $\theta$ is the network parameters. A trajectory, $\tau=(\mathbf{s}_{0},\mathbf{a}_{0},r_{1},\mathbf{s}_{1},\mathbf{a}_{1},r_{2},\mathbf{s}_{2},\mathbf{a}_{2},\ldots,\mathbf{s}_{T})$, is a sequence of trading actions and states observed over a period, where $\mathbf{s}_{0}$ is the initial state and $T$ is the number of steps. The probability of the trajectory $\tau$ is
$P(\tau\mid\pi_{\theta})=\rho_{0}(\mathbf{s}_{0})\prod_{t=0}^{T-1}P(\mathbf{s}_{t+1}\mid\mathbf{s}_{t},\mathbf{a}_{t})\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t}),$ (1)
where $\rho_{0}(s_{0})$ is the initial state distribution. Financial outcomes are quantified as rewards along $\tau$ to compute the cumulative return $R(\tau)=\sum_{t=1}^{T}\gamma^{t-1}r_{t}$. The goal is to learn a policy $\pi_{\theta}$ that maximizes the expected return:
$J(\theta)=\int_{\tau}P(\tau\mid\pi_{\theta})R(\tau)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)],$ (2)
The gradient of $J(\theta)$ with respect to $\theta$ is 27:
@@
where $b\in\mathbb{R}$ is a constant and subtracting $b$ can reduce the variance in practice.
@@
To estimate $\nabla J(\theta)$ in (3), we can use the Monte Carlo method 28:
$\nabla J(\theta)=\frac{1}{n}\sum_{j=1}^{n}\left((R(\tau^{(j)})-b)\sum_{t=0}^{T-1}\penalty 10000\ \nabla_{\theta}\log\pi_{\theta}(\mathbf{a}_{t}^{(j)}\mid\mathbf{s}_{t}^{(j)})\right),$ (4)
@@ 4.1 Automated Data Curation Pipeline for Dynamic Market Data
turbulance Turbulence To control the risk in a worst-case scenario, such as the financial crisis of 2007–2008, FinRL employs the financial turbulence index that measures extreme asset price fluctuation.
@@
Data Sources. We leverage both structured market data and unstructured financial documents. Structured market data captures quantitative market behavior, such as historical price information and market indicator data. Unstructured financial documents provide information about economic trends, regulations, and events that drive market movements. The structured market data are processed into price features $\mathbf{p}_{t}$, which are included in state $\mathbf{s}_{t}$. We outline the financial data utilized in FinRL as follows:
  • OHLCV data. OHLCV (open, high, low, close, volume) data is typical historical volume-price data in finance. We provide daily OHLCV data for stocks from 1999 to 2023. This dataset is prepared through yfinancehttps://yfinance-python.org/ (an open-source Python library) and FinRL-Meta 9, 10, released under the CDLA-permissive-2.0 license. The missing values are removed. We also provide data APIs to download market data, such as Alpacahttps://alpaca.markets/, which are held on the open-source repository FinRL-Meta§§§https://github.com/AI4Finance-Foundation/FinRL-Meta 9, 10.
@@
  • Financial news. The Financial News and Stock Price Integration Dataset (FNSPID) 29 contains 15 million time-aligned financial news articles from 1999 to 2023 for constituent companies in S&P 500. It is released under CC-BY-NA 4.0 for academic purposes. From the FNSPID dataset, we randomly select one news article per day per stock. The news is aggregated with OHLCV data to ensure temporal alignment and create a multimodal dataset. We also provide data APIs to access financial news, such as Yahoo Finance and Finnhub. These data APIs are held on the open-source repository FinGPThttps://github.com/AI4Finance-Foundation/FinGPT 23.
Feature Engineering. Before the dataset can be used in contests, we perform feature engineering.
@@
  • ML-learned factors. For the LOB data, bid and ask prices at different market depths are extracted, and 101 adapted alpha factors 30 are constructed. An RNN is trained to distill 8 strong factors from the 101 weak alphas, modeling and predicting the prices. These factors are used to enrich the feature vector $\mathbf{f}_{t}$ in state $\mathbf{s}_{t}$.
One can perform feature selection as follows: 1) computing Pearson correlations of the features and obtaining a correlation matrix, 2) selecting one feature from a group of correlated ones as in 31.
LLM-Generated Signals. LLMs have been envisioned to be very promising for generating alpha signals from multiple financial data sources 24, 32, 33. We encourage participants to leverage LLMs and train a trading agent for making informed decisions. Here we show the sentiment and risk scores generated by DeepSeek models from financial news 26:
  • Sentiment score. An LLM assigns a sentiment score $u$ of 1 to 5 according to the news, with 1 for strongly negative and 5 highly positive. For example, we can use DeepSeek-V3 34 to generate signals. It is included in the feature vector $\mathbf{f}_{t}$ and is used to adjust actions via the sentiment factor $l^{i}_{t}=1+0.05(u-3)\text{sign}(a^{i}_{t})$. The factor is close to 1 for the stability of the algorithm. The adjusted action is $a_{t}^{i^{\prime}}=l^{i}_{t}a_{t}^{i}$, which is amplified under positive sentiment and dampened under negative sentiment.
  • Risk level. An LLM assigns a risk level $q$ of 1 to 5 from the news, with 1 for low risk and 5 high risk. It is included in the feature vector $\mathbf{f}_{t}$ and is used to penalize rewards through the risk factor $m_{t}^{i}=1+0.05(q_{t}^{i}-3)$. The aggregated risk factor is $M_{t}=\sum_{i}^{K}w_{t}^{i}{m}_{t}^{i}$, where $w_{t}^{i}$ is the portfolio weight of the stock $i$ and $\sum w^{i}=1$. $M_{t}>1$ penalizes the reward for high risk; $M_{t}<1$ increases the reward for low risk.
Automated Data Curation Pipeline. After preparing the datasets, we split them into two parts, one released to participants for training and the other held back for evaluation. We use the temporal partitioning approach to maintain the temporal integrity of financial data and prevent future information leakage:
@@
  • Dataset withheld for evaluation. Depending on the task and data availability, we use two methods to obtain the evaluation dataset: (1) we extract the most recent $\sim$15% of the original dataset, withhold it as the evaluation dataset, and encrypt all timestamps; (2) we collect out-of-sample data for the period after the model submission deadline. It involves downloading market data and financial news via yfinance or other APIs. These two methods will avoid future data leakage and ensure a fair assessment. This setup ensures that the evaluation simulates real-world, time-forward deployment, where models should generalize to new, unseen market data.
@@
4:  Using data period [zYX,zY1][z-Y-X,z-Y-1] to train an agent,
5:  Using data period [zY,z1][z-Y,z-1] to validate the trained agent and adjust hyper-parameters,
@@
7:  Using data period [zYX,z1][z-Y-X,z-1] to retrain the agent with the adjusted hyper-parameters,
@@
9:  Using the trained agent to trade in the $z$-th day.
@@
Alg. 1 summarizes the pipeline of paper trading. For $Z$ trading days (z=0,1,,Z1z=0,1,...,Z-1), we keep doing the following three steps:
  • Step 1). Download and process $X$-day data, from day zYXz-Y-X to day zY1z-Y-1. Then build the data into a gym-style environment and train the agent. Then download and process Y-day data, from day zYz-Y to day z1z-1. Then build the data into a gym-style environment and validate how the agent performs. According to the agent’s performance on the validation environment, adjust hyper-parameters.
  • Step 2). Build the training and validation data, totally $X+Y$ days from day zYXz-Y-X to day z1z-1, into a gym-style environment. Update hyper-parameters to the values chosen from Step 1). Then retrain the agent on these $X+Y$-day environment.
@@
Before each trading day $z$, we use the historical data period from zYXz-Y-X to zY1z-Y-1 to train the FinRL agent. Then, data periods from zYz-Y to z1z-1 are used to validate the trained agent, adjusting hyperparameters accordingly. We retrain the agent by combining all the data from the training and validation period, that is, the data period from zYXz-Y-X to z1z-1. Finally, we perform $1$ day paper trading on the $z$-th day. We continue this process for z=0,1,,Z1z=0,1,...,Z-1.

### v4

@@ 3.2 MDP Formulation
We formulate stock and crypto trading tasks as Markov Decision Processes (MDPs) 6.
@@
  • Action $\mathbf{a_{t}}\in\mathbb{R}^{K}$ represents trading actions at time $t$, such that $\mathbf{h}_{t+1}=\max(\mathbf{h_{t}}+\mathbf{a_{t}},0)$ (making $\mathbf{h}_{t+1}$ nonnegative). An entry $a_{t}^{i}>0,i=1,\ldots,K$ indicates buying $a_{t}^{i}$ shares of assets $i$, while $a_{t}^{i}<0$ indicates selling and $a_{t}^{i}=0$ indicates holding the current position. Each entry of $\mathbf{h_{t+1}}$ is nonnegative, which means the agent cannot sell more shares than it currently holds.
@@
  • Reward $r_{t+1}=R(\mathbf{s_{t}},\mathbf{a_{t}},\mathbf{s_{t+1}})$ is an incentive signal to encourage or discourage the trading agent to perform action $\mathbf{a_{t}}$ at state $\mathbf{s_{t}}$. We use $r_{t+1}$ instead of $r_{t}$ to reflect the delayed reward in financial markets. It emphasizes that the reward $r_{t+1}$ is determined at the next state $\mathbf{s}_{t+1}$ after taking action $\mathbf{a}_{t}$. The reward can be defined as the change in the total asset value, i.e., $R(\mathbf{s_{t}},\mathbf{a_{t}},\mathbf{s_{t+1}})=v_{t+1}-v_{t}$, where $v_{t}=b_{t}+\mathbf{p_{t}}^{T}\mathbf{h_{t}}$ is the total asset value at time $t$. The reward can also be penalized for high-risk signals.
@@
In the state $\mathbf{s}_{t}$, the holding positions $\mathbf{h}_{t}$ are typically constrained to be nonnegative in long-only trading settings, where the agent is not allowed to short assets. The feature vector $\mathbf{f}_{t}$ can include market indicators (e.g., moving average convergence divergence), factors learned by neural networks, and LLM-engineered signals from financial news or regulatory filings. The action $\mathbf{a}_{t}$ can be discrete or continuous, depending on the task. It may support long-only or long-short strategies. The reward function $R$ can use the Sharpe ratio in addition to the change in asset value. It can be further adjusted by risk or market sentiment to better reflect real-world trading scenarios.
For example, a trading task is to develop a trading agent for all 30 constituent stocks in the Dow Jones Index with an initial investment of \$$\$1$ million. In this setting, the number of assets is $K=30$. The state includes the account balance of $b_{0}=100,000$, prices of the 30 stocks, holding positions of the 30 stocks, and $I=4$ technical indicators for each stock. The state space has a dimension $K(I+2)+1=181$. The action space has dimension $K=30$. An entry, for example, $a_{t}^{1}=5$ means buying $5$ shares of the first stock in the portfolio. The reward function is the change in the total asset value, with $v_{0}=100,000$. Taking the first stock as an example, assume at time $t$, the account balance is $b_{t}=100,000$, the price of the first stock is $p_{t}^{1}=200$, and the shares of the first stock are $h_{t}^{1}=10$. The agent buys 5 shares of the first stock ($a_{t}^{1}=5$) and the price increases to $p_{t+1}^{1}=201$ at time $t+1$. Assuming holding all other 29 stocks, the state is updated where $b_{t+1}=99,000$, $p_{t+1}^{1}=201$, and $h_{t+1}=15$. The return $r_{t}=v_{t+1}-v_{t}=(b_{t+1}+\mathbf{p}_{t+1}^{T}\mathbf{h}_{t+1})-(b_{t}+\mathbf{p}_{t}^{T}\mathbf{h}_{t})=b_{t+1}-b_{t}+p_{t+1}^{1}h_{t+1}^{1}-p_{t}^{1}h_{t}^{1}=-1000+201\cdot 15-200\cdot 10=15$. In this example, the agent only operates on the first stock.
@@ 3.3 Policy Gradient Estimate and Its Parallelism
In the FinRL tasks, the policy $\pi_{\theta}$ is typically a neural network (e.g., transformer network or diffusion model), where $\theta$ denotes the learnable parameters. A trajectory, $\tau=(\mathbf{s}_{0},\mathbf{a}_{0},r_{1},\mathbf{s}_{1},\mathbf{a}_{1},r_{2},\mathbf{s}_{2},\mathbf{a}_{2},\ldots,\mathbf{s}_{T})$, is a sequence of trading actions and states observed over a period, where $\mathbf{s}_{0}$ is the initial state and $T$ is the number of steps. The probability of the trajectory $\tau$ is
$P(\tau\mid\pi_{\theta})=\rho_{0}(\mathbf{s}_{0})\prod_{t=0}^{T-1}\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t})P(\mathbf{s}_{t+1}\mid\mathbf{s}_{t},\mathbf{a}_{t}),$ (1)
where $\rho_{0}(s_{0})$ is the initial state distribution. The financial results are quantified as rewards along $\tau$ to calculate the cumulative return $R(\tau)=\sum_{t=1}^{T}\gamma^{t-1}r_{t}$. The goal is to learn a policy $\pi_{\theta}$ that maximizes the following expected return
$J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)]=\int_{\tau}P(\tau\mid\pi_{\theta})R(\tau),$ (2)
The gradient of $J(\theta)$ with respect to $\theta$ is  26
@@
where $b\in\mathbb{R}$ is a constant, and subtracting $b$ can reduce the variance in practice.
@@
To estimate $\nabla J(\theta)$ in (3), we can use the Monte Carlo method 27
$\nabla J(\theta)=\frac{1}{n}\sum_{j=1}^{n}\left((R(\tau^{(j)})-b)\sum_{t=0}^{T-1}\leavevmode\nobreak\ \nabla_{\theta}\log\pi_{\theta}(\mathbf{a}_{t}^{(j)}\mid\mathbf{s}_{t}^{(j)})\right),$ (4)
@@ 4.1 Automated Data Curation Pipeline for Dynamic Market Data
turbulence Turbulence To control the risk in a worst-case scenario, such as the financial crisis of 2007–2008, FinRL employs the financial turbulence index that measures extreme asset price fluctuation.
@@
Data Sources. We leverage both structured market data and unstructured financial documents. Structured market data captures quantitative market behavior, such as historical price information and market indicator data. Unstructured financial documents provide information about economic trends, regulations, and events that drive market movements. The structured market data is processed into price features $\mathbf{p}_{t}$, which are included in state $\mathbf{s}_{t}$. We outline the financial data utilized in FinRL as follows:
  • OHLCV data. OHLCV (open, high, low, close, volume) data is typical historical volume-price data in finance. We provide daily OHLCV data for stocks from 1999 to 2023. This dataset is prepared through yfinancehttps://yfinance-python.org/ (an open-source Python library) and FinRL-Meta 5, 6, released under the CDLA-permissive-2.0 license. The missing values are removed. We also provide data APIs to download market data, such as Alpacahttps://alpaca.markets/, which are held on the open-source repository FinRL-Meta§§§https://github.com/AI4Finance-Foundation/FinRL-Meta 5, 6.
@@
  • Financial news. The Financial News and Stock Price Integration Dataset (FNSPID) 28 contains 15 million time-aligned financial news articles from 1999 to 2023 for constituent companies in S&P 500. It is released under CC-BY-NA 4.0 for academic purposes. From the FNSPID dataset, we randomly select one news article per day per stock. The news is aggregated with OHLCV data to ensure temporal alignment and create a multimodal dataset. For cryptocurrency news, we aggregate BTC news collected from multiple sources The aggregated BTC news dataset: https://huggingface.co/datasets/SecureFinAI-Lab/FinRL_BTC_news_signals. Data sources: https://huggingface.co/datasets/edaschau/bitcoin_news; https://github.com/soheilrahsaz/cryptoNewsDataset and ensure the temporal alignment with the LOB data. We also provide data APIs to access financial news, such as Yahoo Finance and Finnhub. These data APIs are held on the open-source repository FinGPT******https://github.com/AI4Finance-Foundation/FinGPT 22.
Feature Engineering. We conduct feature engineering before making the dataset available for contests.
@@
  • ML-learned factors. For the LOB data, bid and ask prices at different market depths are extracted, and 101 adapted alpha factors 29 are constructed. An RNN is trained to distill 8 strong factors from the 101 weak alphas, modeling and predicting the prices. These factors are used to enrich the feature vector $\mathbf{f}_{t}$ in state $\mathbf{s}_{t}$.
The feature selection process involves two steps: 1) compute the Pearson correlation matrix of the features, and 2) select one representative feature from each group of highly correlated features, following the method proposed by Gort et al 30.
LLM-Engineered Signals. LLMs have been envisioned to be very promising for generating alpha signals from multiple financial data sources 23, 31, 32. We encourage participants to leverage LLMs and train a trading agent for making informed decisions. Here we show the sentiment and risk scores generated by DeepSeek models from financial news 25:
  • Sentiment score. An LLM assigns a sentiment score $u$ of 1 to 5 according to the news, with 1 for strongly negative and 5 for highly positive. For example, we can use DeepSeek-V3 33 to generate signals. It is included in the feature vector $\mathbf{f}_{t}$ and is used to adjust actions via the sentiment factor $l^{i}_{t}=1+0.05(u-3)\text{sign}(a^{i}_{t})$. The factor is close to 1 for the stability of the algorithm. The adjusted action is $a_{t}^{i^{\prime}}=l^{i}_{t}a_{t}^{i}$, which is amplified under positive sentiment and dampened under negative sentiment.
  • Risk level. An LLM assigns a risk level $q$ of 1 to 5 from the news, with 1 for low risk and 5 for high risk. It is included in the feature vector $\mathbf{f}_{t}$ and is used to penalize rewards through the risk factor $m_{t}^{i}=1+0.05(q_{t}^{i}-3)$. The aggregated risk factor is $M_{t}=\sum_{i}^{K}w_{t}^{i}{m}_{t}^{i}$, where $w_{t}^{i}$ is the portfolio weight of the stock $i$ and $\sum w^{i}=1$. $M_{t}>1$ penalizes the reward for high risk; $M_{t}<1$ increases the reward for low risk.
Automated Data Curation Pipeline. After preparing the datasets, we split them into two parts, one released to participants for training and the other held by organizers for evaluation purposes. We use the temporal partitioning approach to maintain the temporal integrity of financial data and prevent future information leakage.
@@
  • Dataset withheld for evaluation. Depending on the task and data availability, we use two methods to obtain the evaluation dataset: 1) we extract the most recent $\sim$15of the original dataset, withhold it as the evaluation dataset, and encrypt all timestamps; 2) we collect out-of-sample data for the period after the model submission deadline. It involves downloading market data and financial news via yfinance or other APIs. These two methods will avoid future data leakage and ensure a fair assessment. This setup ensures that the evaluation simulates real-world, time-forward deployment, where models should generalize to new, unseen market data.
@@
4:     Train agent on data period $[z{-}Y{-}X,\ z{-}Y{-}1]$
5:     Validate trained agent and tune hyperparameters on data period $[z{-}Y,\ z{-}1]$
@@
7:     Retrain agent on data period $[z{-}Y{-}X,\ z{-}1]$ using tuned hyperparameters
@@
9:     Trade on day $z$ with trained agent
@@
Alg. 1 summarizes the pipeline of paper trading. For $Z$ trading days ($z=0,1,...,Z-1$), we keep doing the following three steps:
  • Step 1). Download and process $X$-day data, from day $z-Y-X$ to day $z-Y-1$. Then build the data into a gym-style environment and train the agent. Then download and process Y-day data, from day $z-Y$ to day $z-1$. Then build the data into a gym-style environment and validate how the agent performs. According to the agent’s performance on the validation environment, adjust the hyperparameters.
  • Step 2). Build the training and validation data, totally $X+Y$ days from day $z-Y-X$ to day $z-1$, into a gym-style environment. Update hyperparameters to the values chosen from Step 1). Then retrain the agent on this $X+Y$-day environment.
@@
Before each trading day $z$, we use the historical data period from $z-Y-X$ to $z-Y-1$ to train the FinRL agent. Then, data periods from $z-Y$ to $z-1$ are used to validate the trained agent, adjusting hyperparameters accordingly. We retrain the agent by combining all the data from the training and validation period, that is, the data period from $z-Y-X$ to $z-1$. Finally, we perform $1$ day paper trading on the $z$-th day. We continue this process for $z=0,1,...,Z-1$.